OpenStudy (anonymous):

find the values of the remaining trigonometric functions at t from the given infromation

Parth (parthkohli):

Hmm?

OpenStudy (anonymous):

oops sorry.... for example : sint=5/13 , cos t = -12\13

Parth (parthkohli):

OK. The remaining trig ratios are \(\tan(t), \cot (t), \sec(t), \csc (t)\). First, \(\tan(t) = \dfrac{\sin(t)}{\cos(t)}\)

OpenStudy (anonymous):

oooohhh!! so i would just plug in the numbers into the triangle?

Parth (parthkohli):

Yes, indeed!

Parth (parthkohli):

You don't even need to draw a triangle here. It's straightaway identities. :)

OpenStudy (anonymous):

Ahhh!! See I was just reading the question wrong! :)

Parth (parthkohli):

It's only asking you the values of tan, cot, sec and csc for t.

OpenStudy (anonymous):

What if they give me this : sint= -1\2, cost>0

Parth (parthkohli):

This one is simple as well! You can use the Pythagorean Theorem.

OpenStudy (anonymous):

because sin is -1/2, would cos be the square root of 3/2 ?

Parth (parthkohli):

Yup!

OpenStudy (anonymous):

YOU ARE TO AWESOME MAN!! THANK YOU!! CANT BELIEVE I DIDNT UNDERSTAND THAT! :D Thanks man :)

Parth (parthkohli):

No sweat, bro. Now you do!

OpenStudy (anonymous):

oh one more thing? they will specify a quadrant right? because it could also be -square root of 3/2